OpenStudy (anonymous):

Multiply (3.3 × 104) × (3 × 1018). Express the answer in scientific notation

OpenStudy (anonymous):

10^a * 10^b = 10^(a + b) <=> (3.3 * 10^4) * (3 * 10^18) = (3.3 * 3) * (10^4 * 10^18) = 9.9 * 10^(4 + 18) = 9.9 * 10^22